OpenStudy (anonymous):

What is the graph of the function rule y=3x-2

Vocaloid (vocaloid):

are there any pictures? or does it want you to draw your own?

OpenStudy (anonymous):

I will take a picture

OpenStudy (anonymous):

y=mx+c

OpenStudy (anonymous):

Its a series of graphs with only 1 line graphed, some are vertical, i can't post a picture.

OpenStudy (anonymous):

I tried setting x to 0 and solving for y,i got 1. My niece said shes not sure what they're asking and i don't know without the example

OpenStudy (mathmale):

y=3x-2 is an example of a linear function. The graph is a straight line. The sloe is 3 and the y-intercept is (0,-2). Graph this intercept. Then, find 1 or 2 more points on the graph by starting at (0,-2) and increasing x by 1 and y by 3 (which follows the slope, 3, which is 3/1, or rise/run = 3/1.